OpenStudy (anonymous):

How do i solve -3e^(2X)+16=5? Help pleaseeeeee

OpenStudy (amoodarya):

-3e^(2X)+16=5 -3e^(2X)=5-16 -3e^(2X)=-11 e^(2X)=11/3 now get log( base e ) from both sides

OpenStudy (anonymous):

\[-3e ^{2x}=-11 \rightarrow e ^{2x}=\frac{ 11 }{ 3 }\rightarrow 2x=\ln(\frac{ 11 }{ 3 })\]

OpenStudy (anonymous):

Thanks both of you!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

rm

OpenStudy (anonymous):

NP